OpenStudy (anonymous):

how would i find the limit as x approaches 0 of cos(cos theta) / sec theta

OpenStudy (anonymous):

\[\frac {\cos(\cos \theta)}{( \frac{1}{\cos \theta})}\]

OpenStudy (anonymous):

thank you!